Using the Pythagorean Theorem where a and c are known and b is unknown. How do I solve this problem.
Drive A is heading East and Driver B is heading North. After 30 mins Driver A has gone 20 miles and Driver B has gone 21 miles. How far apart are Drivers A and B after 30 mins of traveling?

Found 2 solutions by davidrolf, Cromlix:
Answer by davidrolf(2)

Hi.
I call distance apart for c
c=sqrt(21^2+20^2)
c=29 miles

Answer by Cromlix(4381)

Hi there,
Consider a right angled triangle.
Driver heading East is the base
Driver heading North is the upright.
Distance apart is the hypotenuse.
A^2 + B^2 = C^2
20^2 + 21^2 = C^2
400 + 441 = C^2
C = √(400 + 441)
C = √(841)
C = 29 miles
Hope this helps :-)